# Tidal phenomena in the Galactic Center: The Curious Case of X7

### v1

@@ 5.2 Ejecta from a stellar merger: are X7 and G3 dynamically linked?
Since X7 and G3 have similar orbits and emission characteristics they have been hypothesized to have been formed in an EKL induced merger scenario where G3 is the merger product and X7 is the ejected mass (Ciurlo et al., 2023). Notice that even the orientations of the angular momentum of X7 and G3 are strikingly similar (see Figure 3). To test this hypothesis, we did a similar analysis as the one done for S33/S0-30 and X7. We extrapolated the orbit of G3 and X7 backwards in time in order to find their closest separation over the last 200 yr period. As a result, we found that this occurred on two occasions: in 1880 and 1952. However, we noticed that G3 is always orbiting ahead of X7 on the sky. Their three-dimensional separation in the closest encounters is larger than 1200 au. Even taking the cases of the closest encounters as the origin of the simulated cloud, we ended up obtaining significant differences between the expected position of the tip of the simulated cloud and X7 in 2021. In the cases of 1880 yr and 1952 yr, we found that the separations are 40000 au and 1200 au between their tips, respectively. When we consider the uncertainties in the orbit of X7, the three-dimensional separation remains of the same order of magnitude in both cases. Moreover, the sky-projected position for the best cases in 1952 are offset beyond the uncertainties as shown in Figure 14.
These results show that ejecta from G3 in 1880 or 1952 cannot be placed on an orbit like the one observed for X7. Thus, it is difficult to reconcile the idea of X7 being ejecta from G3 over the last 200 yr. We note that this conclusion relays on the current orbital estimates of X7 and G3 which have relatively large uncertainties due to short orbital phase coverage (10%; Ciurlo et al., 2023).
@@ 5.3 Grazing collision of a star with a field object
The case of the grazing collision of a star and a field object such as a stellar mass black hole or even a Jupiter-mass object has been discussed in detail by Ciurlo et al. (2023). Such collisions of red giants and compact remnants have also been shown to explain the depletion of red giants within the inner 10″ of the Galactic center (Davies et al., 2011). In principle, a grazing collision of a red giant with any such field objects can account for the mass of X7 by stripping enough material from the star’s atmosphere (Ciurlo et al., 2023). This is an interesting prospect since our simulations show that an initially elongated ejecta from S33/S0-30 with an ejection velocity of 610 can indeed be placed on an orbit similar to X7’s and can reproduce its observed tail orientation (see Figure 8). However, if that were to be the case, the star would be left in an agitated state and would only settle down in the Kelvin-Helmholtz timescale (Ciurlo et al., 2023). There is no account of such effects for S33/S0-30. Moreover, it is hard to explain the existence of other gas and dust sources similar to X7 in the Galactic center, as the collision rate between the red giants and field objects could be as low as once in every years (Rose et al., 2020, 2022, 2023). Albeit being unlikely, our results show a good agreement with this scenario.
@@ 5.4 An infalling gas filament
An additional hypothesis not studied yet is considering X7 as a gaseous clump or stream resulting from the many stellar wind collisions that take place in the vicinity of Sgr A*. Although single stellar wind collisions have been shown to produce only light clumps (0.01 M⊕; Calderón et al., 2020a), in principle the simultaneous interaction of more stellar winds could result in larger and/or more massive structures. Hydrodynamic simulations of the WR stellar winds feeding Sgr A* have been performed by many authors (Cuadra et al., 2005, 2006, 2008, 2015; Russell et al., 2017; Ressler et al., 2018; Calderón et al., 2020b; Ressler et al., 2020; Solanki et al., 2023; Balakrishnan et al., 2024; Calderón et al., 2025). Here, we analyzed the hydrodynamic models developed by Calderón et al. (2025) in order to search for cold and gaseous structures that might share properties with X7. We selected an output of the simulation that corresponds to the last observation of X7, i.e. . Then, we analyzed the region within 10″ (0.4 pc) from Sgr A*. We imposed constraints on gas temperature of 105 K, inclination , and longitude of ascending node in order to find cold structures oriented similarly to that of X7. The results of this analysis are shown as a Hammer projection in Figure 15. Here it is possible to observe that there seem to be cells with similar angular momentum direction to X7’s. However, we checked whether such cells form a coherent spatial structure and found that these are not spatially connected. Thus, it is unlikely that X7 has originated from Wolf-Rayet stellar wind interactions.
@@ 6 Conclusion
We have studied the dynamical evolution of the source X7 in order to constrain its origin. Under the assumption that this object is purely a gas and dust feature, we tested three scenarios for its formation: from the wind of a star going through a high mass-loss episode, as the ejecta of a binary merger, or as the ejecta from the grazing collision of a star with a stellar-mass black hole or Jupiter-mass object. In order to test these hypotheses, we analyzed 195 stars with observationally constrained orbits within the central parsec of the GC. We searched for suitable progenitors with closest approach to X7 of 1000 au, relative velocities of 1000 over the last 200 years, and similar angular momentum orientation. According to these criteria, we found only one suitable candidate: the star S33/S0-30 (Gillessen et al., 2009, 2017; von Fellenberg et al., 2022). This star and X7 had their closest approach in the year 1947 with a minimum separation and relative velocity of 600 au and 500 , respectively. We have modelled the evolution of hypothetical ejecta with a set of test particles launched from the position of S33/S0-30 in 1947 with an initial velocity that minimizes the difference between the three-dimensional distance and relative velocity between the simulated cloud and X7. This analysis shows that:
@@
It is unlikely for X7 to have formed from a stellar wind. No star with sufficiently high mass loss has been close enough to its position over the last 200 years.
@@
Despite their similar orbits X7 and G3 do not seem to be related, as extrapolating their positions to the past does not result in close encounters, and neither can ejecta from G3 form the orbit of X7. It is relevant to bear in mind that this conclusion relies on the age of X7 being less than its orbital period of 200 years.
@@
Initially elongated ejecta of length au from the star S33/S0-30 with an initial velocity of 610 and an initial velocity gradient from tip to tail, can be placed on a similar orbit to X7’s, and reproduce the current position of its tip. Additionally, the length of such ejecta is 3350 au in 2021 which is consistent with the observational data.
@@
Secondary effects such as stellar wind ram pressure, and hypothetical feedback event from the SMBH are not significant to affect the evolution of X7. A static ISM drag can indeed change the orbit of X7 but it does not affect its dynamics significantly over the simulated timescale of 80 yr.
From these results, we speculate that a grazing collision between S33/S0-30 and a field object such as a stellar-mass black hole or a Jupiter-mass object could have produced the ejecta that corresponds to X7. It will be interesting to observe the star S33/S0-30 for any extended dusty envelope which could further strengthen our results. As X7 approaches the pericenter passage, it will elongate more and begin fragmenting, which has already been observed. Our simulations are fairly accurate before the pericenter passage. To reveal more about the tidal interaction and post-pericenter evolution, numerical hydrodynamic simulations are required. Future observations will reveal more interesting properties with better constraints.
@@ Acknowledgements.
WS and JC acknowledge the kind hospitality of the Max Planck Institute for Astrophysics where part of this work was developed. The research of WS, DC, and SR has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 - “Quantum Universe” - 390833306. Since 01.10.2024, DC has been funded by the Alexander von Humboldt Foundation. SR has also been supported by the Swedish Research Council (VR) under grant number 2020-05044, by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012, by the Knut and Alice Wallenberg Foundation under grant Dnr. KAW 2019.0112, and by the European Research Council (ERC) Advanced Grant INSPIRATION under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101053985). JC acknowledges financial support from ANID – FONDECYT Regular 1211429 and 1251444, and Millennium Science Initiative Program NCN. The analysis and visualizations in this work made use of Python libraries, Numpy (Harris et al., 2020) and Matplotlib (Hunter, 2007). RAMSES hydrodynamical simulation data analysis was carried out using the Python library YT (Turk et al., 2011).Data Availability
@@ Appendix A Spherical initial configuration
In the main text, we presented the result of simulations that considered an initially ellipsoidal cloud as that showed the best agreement with the observations. For completeness, here we included the results considering a uniform initially spherical cloud. This simulation resulted in a three-dimensional separation between the simulated cloud and the observed tips of X7 of au in 2021. The length of the simulated cloud was 2000 au in 2021 whereas the observations show a length of 3300 au (Ciurlo et al. 2023). However, the main discrepancy comes from the orientation of the tail of the simulated cloud and X7. Figure 16 shows the sky-projected morphology of the simulated cloud at and yr as dark blue, light blue, cyan, dark green, and light green dots, respectively. The gray shaded lines indicate the orientation of the observed structure of X7 (Ciurlo et al. 2023). The dashed red lines are lines connecting Sgr A* and the observed positions of the tip of X7 at the reported times. The location of Sgr A* is on the uppermost left-hand side corner. The disagreement between the orientation of the simulated and observed structures is clear. The discrepancy is around during the epoch 2002-2021 yr.
@@ Appendix B A different orbit for X7
Peißker et al. (2024b) derived a different orbit for X7 when compared to the one by Ciurlo et al. (2023). A comparison between these orbits can be seen on the left-hand side panel of Figure 17. The solid red and black lines represent the orbits inferred by Ciurlo et al. (2023) and Peißker et al. (2024b), respectively. It is clear that there is a large discrepancy between the two orbits. Although the uncertainties in Peißker et al. (2024b) are smaller by an order of magnitude, the key reason for this different orbit is their L′-band data point from 1999 which favors this newly derived orbit.
In order to test if the results of our work are affected by the X7 orbit choice we repeated the analysis but making use of the orbit derived by Peißker et al. (2024b). The calculations showed that the closest stars to X7 within the last 200 years are at least 4000 au apart (see right-hand side panel of Figure 17). Bear in mind that our original separation threshold was 1000 au. As in this case the separation is at least four times larger the use of a different orbit makes less likely the relation to any of the star candidates.
@@ Appendix C Simulated morphology evolution
For completeness, we analyzed all the Br- observations available and compared them with the simulated emission. Figure 18 shows the sky-projected Br- flux observed from Earth through the epochs , 2008, 2009, 2010, 2011, 2012, 2013, 2014, 2017, 2018, 2019, and 2021. Each panel corresponds to 1″ 1″with Sgr A* located on the uppermost left-hand side corner. The simulated Br- flux is shown as colormap whereas the observed emission is depicted as colored contours. Notice that the observations contain most of the distribution of the simulated flux, and the emission level is of the same order as the reported values as it is shown in the main text.

### v2

@@ 5.2 Ejecta from a stellar merger: Are X7 and G3 dynamically linked?
Since X7 and G3 have similar orbits and emission characteristics they have been hypothesized to have been formed in an EKL-induced merger scenario where G3 is the merger product and X7 is the ejected mass (Ciurlo et al., 2023). Notice that even the orientations of the angular momentum of X7 and G3 are strikingly similar (see Fig. 3). To test this hypothesis, we conducted an analysis similar to that done for S33/S0-30 and X7. We extrapolated the orbit of G3 and X7 backward in time in order to find their closest separation over the last 200 yr period. As a result, we find that this occurred on two occasions: in 1880 and 1952. However, we notice that G3 is always orbiting ahead of X7 on the sky. Their three-dimensional separation in the closest encounters is larger than 1200 au. Even taking the cases of the closest encounters as the origin of the simulated cloud, we ended up obtaining significant differences between the expected position of the tip of the simulated cloud and X7 in 2021. In the cases of 1880 yr and 1952 yr, we find that the separations are 40000 au and 1200 au between their tips, respectively. When we consider the uncertainties in the orbit of X7, the three-dimensional separation remains of the same order of magnitude in both cases. Moreover, the sky-projected position for the best cases in 1952 are offset beyond the uncertainties as shown in Fig. 14.
These results show that ejecta from G3 in 1880 or 1952 cannot be placed on an orbit like the one observed for X7. Thus, it is difficult to reconcile the idea of X7 being ejecta from G3 over the last 200 yr. We note that this conclusion relies on the current orbital estimates of X7 and G3, which have relatively large uncertainties due to short orbital phase coverage (10%; Ciurlo et al., 2023).
@@ 5.3 Grazing collision of a star with a field object
The case of the grazing collision of a star and a field object such as a stellar-mass black hole or even a Jupiter-mass object has been discussed in detail by Ciurlo et al. (2023). The collisions between red giants and compact remnants have also been invoked to explain the depletion of red giants within the inner 10″ of the GC (Davies et al., 2011). In principle, a grazing collision of a red giant with any such field object could account for the mass of X7 by stripping enough material from the star’s atmosphere (Ciurlo et al., 2023). This is an interesting prospect since our simulations show that an initially elongated ejecta from S33/S0-30 with an initial velocity of 610 can indeed be placed on an orbit similar to X7’s, as well as reproduce its observed tail orientation (see Fig. 8). However, if that were the case, the star would be left in an agitated state and would only settle down on a Kelvin-Helmholtz timescale (Ciurlo et al., 2023). To date, there is no account of such effects for S33/S0-30, but future observations might be able to detect this.
To assess how feasible this scenario is, we need to ask how often (grazing) collisions happen in the GC. There have been studies dedicated to estimating the stellar collision rate, motivated by the missing red giant problem (Rose et al., 2020, 2022, 2023; Amaro Seoane, 2023). These studies estimated this rate to be about yr-1 in a volume of 100 Mpc for a Milky Way like nucleus. Balberg & Yassur (2023) developed a probabilistic approach to study the rates of collisions in a stellar cluster with a SMBH at its center. This work predicted collision rates in the range 10-6- yr-1, smaller, as they referred to destructive, i.e., physical collisions as opposed to just stripping the outer layers of the stars. However, none of these works considered the potential collisions between a field object such as a stellar-mass black hole and a star. The central parsec is thought to host a population of compact objects, including stellar-mass black holes (Morris, 1993; Mouawad et al., 2005; Zhao et al., 2022). Recent work by Haas et al. (2025) calculated the radial density profile of the black hole cluster in the GC based on observations of the stellar populations. Their estimations suggest that at a distance of 0.015 pc from Sgr A*, which is approximately the location of X7, a star of 10 M⊙ should have gone through a collision with a black hole on a timescale of 55 Myr, while a more massive star of 50 M⊙ would do so on a shorter timescale of 5 Myr. Moreover, it is indeed feasible that such collisions have already taken place for some of the stars that inhabit the region, as the young population age is on the order of 2.5-5.8 Myr (Lu et al., 2013). Under this scenario, let us consider the star S33/S0-30 and assume its mass to be 10 M⊙. Currently, it is possible to observe 100 B-type stars like it in the region. Then, we would expect one collision between a B-type star and a stellar-mass black hole every 0.05 Myr. If we consider that a stripped filament, such as X7, lives for 100 yr, the chances of detecting it would be 0.2%. Notice that this value should be interpreted as a lower limit since S33/S0-30 could have been more massive in the past and would have lost mass due to this kind of interaction. Furthermore, we note that the calculations by Haas et al. (2025) do not consider a preferential geometry. In the case of the young stars in the GC, the collision timescale could be shorter if we weigh in the fact that a significant fraction of the stars have been formed on disk structures (Paumard et al., 2006; Gillessen et al., 2017; von Fellenberg et al., 2022). Moreover, it is to be noted that the grazing collision timescales should be even shorter than those computed by Haas et al. (2025) for more direct collisions, given the larger cross section for grazing collisions. Thus, although uncertain, the collision rate estimations between a field object and a star are compatible with X7 having formed in one such event.
@@ 5.4 An infalling gas filament
An additional hypothesis not studied yet is considering X7 as a gaseous clump or stream resulting from the many stellar wind collisions that take place in the vicinity of Sgr A*. Although single stellar wind collisions have been shown to produce only light clumps (0.01 M⊕; Calderón et al., 2020a), in principle the simultaneous interaction of more stellar winds could result in larger and/or more massive structures. Hydrodynamic simulations of the WR stellar winds feeding Sgr A* have been performed by many authors (Cuadra et al., 2005, 2006, 2008, 2015; Russell et al., 2017; Ressler et al., 2018; Calderón et al., 2020b; Ressler et al., 2020; Solanki et al., 2023; Balakrishnan et al., 2024; Calderón et al., 2025). Here, we analyzed the hydrodynamic models developed by Calderón et al. (2025) in order to search for cold and gaseous structures that might share properties with X7. We selected an output of the simulation that corresponds to the last observation of X7, i.e., . Then, we analyzed the region within 10″ (0.4 pc) from Sgr A*. We imposed constraints on gas temperature of 105 K, inclination , and longitude of ascending node in order to find cold structures oriented similarly to that of X7. The results of this analysis are shown as a Hammer projection in Fig. 15. Here it is possible to observe that there seem to be cells with similar angular momentum direction to X7’s. However, we checked whether such cells form a coherent spatial structure and found that these are not spatially connected. Thus, it is unlikely that X7 has originated from WR stellar wind interactions.
@@ 6 Conclusion
We have studied the dynamical evolution of the source X7 in order to constrain its origin. Under the assumption that this object is purely a gas and dust feature, we tested three scenarios for its formation: from the wind of a star going through a major mass-loss episode, as the ejecta of a binary merger, or as the ejecta from the grazing collision of a star with a stellar-mass black hole or Jupiter-mass object. To test these hypotheses, we analyzed 195 stars with observationally constrained orbits within the central parsec of the GC. We searched for suitable progenitors with a closest approach to X7 of 1000 au, relative velocities of 1000 over the last 200 years, and a similar angular momentum orientation. With these criteria, we found only one suitable candidate: the star S33/S0-30 (Gillessen et al., 2009, 2017; von Fellenberg et al., 2022). This star and X7 had their closest approach in the year 1947 with a minimum separation and relative velocities of 600 au and 500 , respectively. We modeled the evolution of hypothetical ejecta with a set of test particles launched from the position of S33/S0-30 in 1947 with an initial velocity that minimizes the difference between the three-dimensional distance and the relative velocity between the simulated cloud and X7. This analysis shows that:
@@
It is unlikely that X7 formed from a stellar wind. No star with sufficiently high mass loss has been close enough to its position over the last 200 years.
@@
Despite their similar orbits, X7 and G3 do not seem to be related: extrapolating their positions into the past does not result in close encounters, and ejecta from G3 cannot form the orbit of X7. It is relevant to bear in mind that this conclusion relies on the age of X7 being less than its orbital period of 200 years.
@@
Initially elongated ejecta of length au from the star S33/S0-30 with an initial velocity of 610 and an initial velocity gradient from tip to tail can be placed on a similar orbit to X7’s, and can reproduce the current position of its tip. Additionally, the length of such ejecta is 3350 au in 2021, which is consistent with the observational data.
@@
Secondary effects such as stellar wind ram pressure and a hypothetical feedback event from the SMBH are not significant enough to affect the evolution of X7. A static interstellar medium drag can indeed change the orbit of X7, but it does not affect its dynamics significantly over the simulated timescale of 80 yr.
From these results, we speculate that a grazing collision between S33/S0-30 and a field object such as a stellar-mass black hole or a Jupiter-mass object could have produced the ejecta that corresponds to X7. Furthermore, we have argued that the current constraints on the presence of a population of stellar-mass black holes in the vicinity of Sgr A* make grazing collisions a plausible occurrence that could account for the production of X7.
It will be interesting to observe the star S33/S0-30 for any extended dusty envelope, which would further strengthen our results. As X7 approaches the pericenter passage, it will elongate more and begin fragmenting, which has already been observed. Our simulations are fairly accurate before the pericenter passage. To reveal more about the tidal interaction and post-pericenter evolution, numerical hydrodynamic simulations are required. Future observations will reveal more interesting properties with better constraints.
@@ Acknowledgements.
We would like to thank the anonymous referee for the comments that helped to improve this article. WS and JC acknowledge the kind hospitality of the Max Planck Institute for Astrophysics where part of this work was developed. The research of WS, DC, and SR has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 - “Quantum Universe” - 390833306. Since 01.10.2024, DC has been funded by the Alexander von Humboldt Foundation. SR has also been supported by the Swedish Research Council (VR) under grant number 2020-05044, by the research environment grant “Gravitational Radiation and Electromagnetic Astrophysical Transients” (GREAT) funded by the Swedish Research Council (VR) under Dnr 2016-06012, by the Knut and Alice Wallenberg Foundation under grant Dnr. KAW 2019.0112, and by the European Research Council (ERC) Advanced Grant INSPIRATION under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101053985). JC acknowledges financial support from ANID – FONDECYT Regular 1211429 and 1251444, and Millennium Science Initiative Program NCN. The analysis and visualizations in this work made use of Python libraries, Numpy (Harris et al., 2020) and Matplotlib (Hunter, 2007). RAMSES hydrodynamical simulation data analysis was carried out using the Python library YT (Turk et al., 2011).Data availability
@@ Appendix A Spherical initial configuration
In the main text, we have presented the result of simulations that considered an initially ellipsoidal cloud as that showed the best agreement with the observations. For completeness, here we include the results considering a uniform initially spherical cloud. This simulation resulted in a three-dimensional separation between the simulated cloud and the observed tips of X7 of au in 2021. The length of the simulated cloud was 2000 au in 2021 whereas the observations show a length of 3300 au (Ciurlo et al. 2023). However, the main discrepancy comes from the orientation of the tail of the simulated cloud and X7. Figure 16 shows the sky-projected morphology of the simulated cloud at and yr and the orientation of the observed structure of X7 (Ciurlo et al. 2023). The disagreement between the orientation of the simulated and observed structures is clear. The discrepancy is around during the epoch 2002-2021 yr.
@@ Appendix B A different orbit for X7
Peißker et al. (2024b) derived a different orbit for X7 when compared to the one by Ciurlo et al. (2023). A comparison between these orbits can be seen on the top panel of Fig. 17. It is clear that there is a large discrepancy between the orbits inferred by Ciurlo et al. (2023) and Peißker et al. (2024b). Although the uncertainties in Peißker et al. (2024b) are smaller by an order of magnitude, the key reason for this different orbit is their L′-band data point from 1999, which favors this newly derived orbit.
To test if the results of our work are affected by the X7 orbit choice, we repeated the analysis but making use of the orbit derived by Peißker et al. (2024b). The calculations show that the closest stars to X7 within the last 200 years are at least 4000 au apart (see the bottom panel of Fig. 17). Bear in mind that our original separation threshold was 1000 au. As in this case the separation is at least four times larger, the use of a different orbit makes a relation to any of the star candidates less likely.
@@ Appendix C Simulated morphology evolution
For completeness, we analyzed all the Br observations available and compared them with the simulated emission. Figure 18 shows the sky-projected Br flux observed from Earth through the epochs , 2008, 2009, 2010, 2011, 2012, 2013, 2014, 2017, 2018, 2019, and 2021. Notice that the observations contain most of the distribution of the simulated flux, and the emission level is of the same order as the reported values as it is shown in the main text.